OpenStudy (anonymous):

Your family is planning a 4 hour car trip. Along the way you are planning to make three 1/2 hour stops. At what time should you leave home to arrive at the destination by 8:00 pm

OpenStudy (anonymous):

1/2 hour = 30 minutes, 60 minutes = 1 hour. So: \[\frac{ 1 }{ 2 } \times 3 = 1 \frac{ 1 }{ 2 } hours\] Add that on to 4 hours: \[4 hours + 1\frac{ 1 }{ 2 } hours = 5 \frac{ 1 }{ 2 } hours\]

OpenStudy (anonymous):

@mangorox so 2:30

OpenStudy (anonymous):

Well you need to work out the total time the trip will take. That means the 4 hours on the road + 3 1/2 hours stops. The time they need to leave at is then 8-TOTALTIME